OpenStudy (anonymous):

If 6x - 3y = 30 and 4x = 2 - y then find x + y.

OpenStudy (anonymous):

i'll do it in a minute

OpenStudy (anonymous):

damn, my system ruined ;(((

OpenStudy (anonymous):

linear system finding the unknown variables. .. to find x and y, here are the ways: 1. substitution method 2.elimination method 3.graphing method 4.discriminant after it, since the question is find the value of (x+y), after getting the value of x and y using the above ways, add the value of it. In that way you gonna find the value of x+ y.

OpenStudy (vishal_kothari):

x+y= -4

OpenStudy (anonymous):

6x - 3y = 30 and 4x = 2 - y then find x + y. 4x-2= -y ......... y= -4x +2 .... 6x-3(-4x+2)=30 .......... 6x+12x -6 = 30 ............. 18x - 6 = 30 ............... 18x=36 ............. x= 2 ........... 4x = 2 - y ...... 4(2) = 2 -y ...... 8 = 2 - y ....... 6 = -y ..... y=-6 ....... x+y= -6 + 2 = -4